Dawit plans to consume two goods (good x and good y) with a limited budget. If Dawit’s budget line has intercepts 100 units of good x and 50 units of good y and price of good x (Px) is $100. Then, answer the next three questions. A) What are Dawit’s income and Price of good y (Py) B) What is the simplified version of Dawit’s budget line equation? C) If Dawit’s utility function is given by U(X; Y) = X0.5Y 0.5, how many units of good X and good Y would he consume if he choose the bundle that maximize his utility subject to his budget constraint?


1
Expert's answer
2021-09-13T12:03:36-0400
"I=100\times100=10,000"

"P_Y=\frac{I}{Q_Y}=200"

"MU_X=0.5(\frac{y}{x})^{0.5}"

"MU_Y=0.5(\frac{x}{y})^{0.5}"

"\frac{MU_X}{p_X}=\frac{MU_Y}{p_Y}"

"x=2y"

"100x+200y=10,000"

"y=25"

"x=50"